The electric field (E) at 1cm away from a point charge is 900 N/C . What is the magnitude of the charge?

a.   1.0 x10-11C

b.   2.7 x10-11C

c.   9.0 x10-9C

d.   2.7 x10-9C
Answer #1

We know that electric field due to a point charge at a distance of r = kq/r^2

So , 900 = kq/r^2

900 = (9×10^9×q)/(0.01)^2

From here we get q = 1×10^(-11)C which is option A